Please help me solve-
How much should be invested now at (i.e. fid the present value) 4% simple interest to have $4620 in 8 years?

Answer by checkley77(12844)

4620=P(1+R)^T
4620=P(1+.04)^8
4620=P(1.04)^8
4620=P*1.3686
P=4620/1.3686
P=$3,375.71 IS THE ORIGINAL AMOUNT INVESTED.